OpenStudy (timinthor):

Which answer is an equation in point-slope form for the given point and slope? Point: (7, 4); Slope: 6 y−4=6(x−7) y−7=6(x+4) y+4=6(x−7) y−4=6(x+7)

OpenStudy (studygurl14):

Point slope form: \(\Large y-y_1=m(x-x_1)\) where m = slope and \((x_1,y_1)\) is a point on the line

OpenStudy (timinthor):

So the answer would be A.? @StudyGurl14

OpenStudy (timinthor):

Yup, it is. I just submitted the test, 100%. Thanks!!

OpenStudy (studygurl14):

yep :)